Nouabalé Ndoki (04.01.2014 to 06.30.2014)

You are here

PDF versionPDF version
Year: 
2014
Quarter: 
April - June
Communications activities: 

None

Unusual Events at the TEAM Site: 

None

New Species: 

None

Protocol Activities: 

 

 

 

VG data for the Array 1 and Array 2 were uploaded on the TEAM portal. Trees and lianas of six plots of 1ha each  were re-measured with a total of 2240 stems (1063 in Array 1 and 1177 in Array 2) and 17 new recruits . A  number of 29 trees/lianas stems changed their POM due to trunk deformation.

 

CL data are uploaded on the TEAM portal on a regular basis. Since the climate station system was updated in January it is functioning normally

 

A  number total of 30 cameras were deployed the last trimester (in Array 1) but data for the 60 cameras are still being processed because after collecting the camera from the field the technician and myself were absent from the site. We were attending to different workshops and seminars (different from the TEAM subject) in and out the country. Data will be ready in the mid-July and once ready, they will be sent on a USB key to WCS HQ.

Regarding the CT for the two events deployments (Array 1 and Array 2) we lost 5 cameras due to termites and elephants activities.

 

 

 

Protocol Problems: 

None

Schedule Problems: 

None

Logistical Problems: 

None

Suggestions: 

It would be beneficial for the field team to have hand held tablets to collect data - particularly for the VG data. This would not only save time but would avoid errors due to data entry. Data could be saved on SD memory card and then uploaded from a laptop directly into the TEAM portal after the field mission